# Channel-Selective Normalization for Label-Shift Robust Test-Time Adaptation

### v1

@@ Abstract
Deep neural networks have useful applications in many different tasks, however their performance can be severely affected by changes in the data distribution. For example, in the biomedical field, their performance can be affected by changes in the data (different machines, populations) between training and test datasets. To ensure robustness and generalization to real-world scenarios, test-time adaptation has been recently studied as an approach to adjust models to a new data distribution during inference. Test-time batch normalization is a simple and popular method that achieved compelling performance on domain shift benchmarks. It is implemented by recalculating batch normalization statistics on test batches. Prior work has focused on analysis with test data that has the same label distribution as the training data. However, in many practical applications this technique is vulnerable to label distribution shifts, sometimes producing catastrophic failure. This presents a risk in applying test time adaptation methods in deployment. We propose to tackle this challenge by only selectively adapting channels in a deep network, minimizing drastic adaptation that is sensitive to label shifts. Our selection scheme is based on two principles that we empirically motivate: (1) later layers of networks are more sensitive to label shift (2) individual features can be sensitive to specific classes. We apply the proposed technique to three classification tasks, including CIFAR10-C, Imagenet-C, and diagnosis of fatty liver, where we explore both covariate and label distribution shifts. We find that our method allows to bring the benefits of TTA while significantly reducing the risk of failure common in other methods, while being robust to choice in hyperparameters.
@@ 1 Introduction
Deep learning models have shown excellent performance on a wide variety of tasks in computer vision research, such as image classification [22, 50], segmentation [46, 11], and object detection [7, 48]. However, a commonly cited limitation of these models is the inability to generalize across different domains [14]. Generalization can be simply defined as the ability of a deep learning algorithm, trained on one or more source domains, to be applied to a different yet still related target domain. Typically, in real-world deployment scenarios models might encounter data with critical differences, hampering their performance. This decrease in performance has been observed in multiple areas, including life-threatening contexts, such as autonomous driving [19, 43] and medical diagnostics [3, 45].
@@
Domain adaptation techniques aim to enhance a model’s generalizability, anticipating the possible challenges in the event of distribution shifts. Traditionally, domain adaptation assumes that the model has access to unlabeled test data and source data during the training, which may not be practical in some cases [34, 25]. A recently emerging technique to deal with distribution shift is test-time adaptation (TTA) [40, 26], a type of unsupervised domain adaptation, where unlabeled test data is used to update the model parameters at test-time, before predictions. It is often assumed that data arrives in batches, and some studies have proposed a setting of test-time batch adaptation that take advantage of batch-level information to adapt to the distribution shift [44, 30, 8].
@@
Test-time batch normalization (TTN) [30, 37] is an approach which replaces batch normalization statistics estimated as running averages on the training set with the statistics of the test data batch. Despite being a simple approach, it has been shown to improve robustness under covariate shift, handling particularly well various cases of image corruptions. Based on that, other TTA approaches apply TTN as a critical component in their foundation [44, 27]. Alas, most existing TTA methods consider the impact of covariate shifts only, in many realistic scenarios, such as medical imaging, the label distribution of data can shift from training to testing.
@@
In order to be practically useful, TTA methods relying on batch data need to be able to operate outside of this optimistic scenario of class-balanced training and testing settings, while simultaneously dealing with covariate shifts. As noted in recent work [4] label distribution shifts can cause failures. We emphasize in our work that sometimes these failures can be catastrophic, thus creating risk in the usage of TTA in the first place when the label distribution of batches obtained under deployment cannot be adequately guaranteed. We observe in our work that certain channels and layers can be much more sensitive with respect to label distribution shifts under TTN. Motivated by this, we propose a method to mitigate the risk of label distribution shift catastrophically affecting model predictions. It is based on the adaptation of only those channels of the batch normalization layers whose activation statistics do not exhibit strong correlation to label distribution. Our proposed method is applied to classification of two well-known benchmark natural image datasets (CIFAR-10 [21] and ImageNet-1K [9]), and finally on adaptation from a private medical dataset [42] to a publicly available one [6]. When deployed in target data with different distribution, our proposed method is effective in obtaining the benefits of TTA while avoiding catastrophic failure cases, greatly improving the practical utility of TTN methods.
@@
    Experiments on three different datasets demonstrate that our method can be applied on a range of classification tasks, in particular for highly imbalanced target data. Notably, the adaptation across liver ultrasound datasets from different sites is shown to be effective.
@@ 2 Related Work
Batch normalization [17], henceforth BatchNorm, is a technique that aims to stabilize the distributions of layer inputs by introducing additional network layers that control the first two moments (mean and variance) of these distributions. This positively affects the optimization process of deep neural networks during training, reduces overfitting, and prevents the exploding or the vanishing gradient problems.
@@
Even though BatchNorm is effective, widely accepted, and commonly used in many deep neural networks, its mechanisms are not entirely understood. Methods extending this approach provide alternatives to BatchNorm, such as layer normalization [1], instance normalization [41], and normalization based on groups of channels [47]. Furthermore, weight normalization [35] extends this line of work, modifying the weights of the models instead of features.
@@
Various studies have been conducted to further understand how batch-normalization works. The original motivation for BatchNorm refers to the problem of internal covariate shift, however a later study [36] notes that it is not an appropriate explanation. Rather, the loss landscapes of models with BatchNorm appear to be smoother than those of equivalent models without it, allowing for use of larger range of learning rates and faster network convergence. Moreover, by investigating randomly initialized networks with all layers frozen and adapting only the BatchNorm layers, it has been shown that BatchNorm layers have substantial expressive power, achieving surprisingly good results on CIFAR-10 and ImageNet [12].
@@
Test-time adaptation methods often involve updating the parameters associated with the BatchNorm layers to distributions with covariate shift [37, 30]. This approach is closely related to the unsupervised domain adaptation method AdaBN [24, 5], and has achieved promising results in mitigating image corruptions. Recent TTA methods have been published using entropy minimization on the BatchNorm affine parameters [44], but they appear to be biased towards the dominant classes [33]. In contrast, a distinct class of methods focuses on TTA without relying on BatchNorm adjustments, proposing techniques that leverage ideas from self-supervised learning [39, 8], or test-time augmentations [16, 18].
@@
A common limitation in TTA studies is regarding label distribution shifts, as most of the proposed methods fail to consider the varying class balance in the test sets, eventually falling into the “winner-takes-all” dilemma in which the majority classes dominate the predictions [49, 26]. To address the class-imbalanced domain adaptation challenge, some works focus on estimating the test set distribution [38, 28, 33].
@@
Recently, other works have also proposed methods for dealing with label distribution shift in TTA. LAME [4] performs the online adaptation of the classification outputs instead of the model’s parameters, making it agnostic to training and testing conditions, but it does not perform well in balanced scenarios. DELTA [53] uses test-time batch renormalization and dynamic online reweighting in an online manner to avoid mini-batch bias. A label shift adapter was designed [33] to improve TTA performance in scenarios where both source and target are imbalanced, by estimating the distribution of the target set. However, it uses a large number of hyperparameters and is specialized to the case of imbalanced source data.
@@ 3 Hybrid Test-Time Batch Normalization
In order to implement TTA in these settings, TTN views a neural network $f$ as split into blocks separated by BatchNorm layers:
$f=f_{0}\circ B_{0}^{s_{0}}\circ f_{1}\circ B_{1}^{s_{1}}\cdots f_{K-1}\circ B_{K-1}^{s_{K-1}}\circ f_{K},$ (1)
where $f_{0},\ldots,f_{K}$ are blocks (i.e., sub-networks) of hidden layers and $B_{0},\ldots,B_{K-1}$ are batch normalization operators. Each BatchNorm layer modifies each neural activation by
@@
Input: Trained network $n$ with $K$ layers; source model statistics $\{\bm{\mu}_{k}^{s},\bm{\sigma}_{k}^{s}\}_{k=1..K}$; source data $x_{s}$;
1 for $c$ in $C$ classes do
2       Sample Data from Class $c$
3       $x_{0}\sim D_{c}$
4       for $k$ in $\{1,...,K\}$ do
5             $\bm{x}_{k}=n_{k-1}(\bm{x}_{k-1})$
6             $\bm{\mu}_{k}^{t},\bm{\sigma}_{k}^{t}$leftarrow Compute BN stats from  $\bm{x}_{k}$
7             for $f$ in $F$ channels do
8                   Scores[c, k, f] = $W_{2}^{2}(\bm{\mu}_{f}^{s},\bm{\sigma}_{f}^{s},\bm{\mu}_{f}^{t},\bm{\sigma}_{f}^{t})$
9
10             end for
11
12       end for
13
14 end for
Output:Scores
Algorithm 1 Hybrid-TTN Post-Training
1 for $k$ in $\{1,...,K\}$ do
2       $\bm{x}_{k}=f_{k-1}(\bm{x}_{k-1})$
3       $\bm{\mu}_{k}^{t},\bm{\sigma}_{k}^{t}$leftarrow Compute BN stats from  $\bm{x}_{k}$
4       $R=\frac{k}{K}$
5       for $c$ in $C$ classes do
6             for $f$ in $F$ channels do
[…]
8             end for
9
10       end for
[…]
13       $\bm{m}_{k}[\tilde{\textsc{topR-ind}}]=1$
14       $\bm{\mu}_{k}^{hybrid}=\bm{m}_{k}\odot\bm{\mu}_{k}^{t}+(1-\bm{m}_{k})\odot\bm{\mu}_{k}^{s}$
15       $\bm{\sigma}_{k}^{hybrid}=\bm{m}_{k}\odot\bm{\sigma}_{k}^{t}+(1-\bm{m}_{k})\odot\bm{\sigma}_{k}^{s}$
16       $\bm{x}_{k}=B(\bm{x}_{k},\bm{\mu}_{k}^{hybrid},\bm{\sigma}_{k}^{hybrid})$
17
18 end for
[…]
Algorithm 2 Hybrid-TTN Adaptation
@@ 16       $\bm{x}_{k}=B(\bm{x}_{k},\bm{\mu}_{k}^{hybrid},\bm{\sigma}_{k}^{hybrid})$
Storing Post-Training Statistics. To mitigate the adverse effects by TTN, our first step focuses on selecting channels for the adaptation process. We aim to identify channels that are primarily sensitive to covariate shifts while excluding those highly sensitive to shifts in the label distribution. This channel selection procedure operates using an initial phase that uses the labelled training set before the model is deployed to compute a set of scores. For each class, denoted as $c$ out of $C$ total classes, we input a subset of data from the class, $D_{c}$, into a trained neural network with $K$ layers. Each layer $k$ within this network possesses source statistics $\mu_{k}^{s}$ and $\sigma_{k}^{s}$, and the target statistics for the single class, $\mu_{k}^{t}$ and $\sigma_{k}^{t}$, are computed. The Wasserstein distance is then calculated for the source and target distributions with $W_{2}^{2}(\{\mu^{s},\sigma^{s}\},\{\mu^{t},\sigma^{t}\})=\|\mu^{s}-\mu^{t}\|^{2}+(\sigma^{s}-\sigma^{t})^{2}$. This step is described in Algorithm 1.
@@
At each layer we choose the top-$R$ channels for adaptation. Using the prior that later layers are more sensitive to label shifts we decay $R$ with a simple linear function in depth. The algorithm is fully described in Algorithm 2.
Re-scoring with Pseudolabels In the method analyzed in our paper we focus on a simple method to obtain $p(c)$. Specifically we first run Algorithm 2 with a uniform prior to obtain an initial corrected model, then use this to obtain pseudolabels for the batch, which are used to recalculate the prior, we then run Algorithm 2 again with the non-uniform prior.

### v2

@@ Abstract
Deep neural networks have useful applications in many different tasks, however their performance can be severely affected by changes in the input distribution. For example, in the biomedical field, their performance can be affected by changes in the data (different machines, populations) between training and test datasets. To ensure robustness and generalization to real-world scenarios, test-time adaptation has been recently studied as an approach to adjust models to a new data distribution during inference. Test-time batch normalization is a simple and popular method that achieved compelling performance on domain shift benchmarks. It is implemented by recalculating batch normalization statistics on test batches. Although initial work focused on analysis with test data that has the same label distribution as the training data, in many practical applications this technique is vulnerable to label distribution shifts, sometimes causing catastrophic failure. This presents a risk in applying test time adaptation methods in deployment. We propose to tackle this challenge by only selectively adapting channels in a deep network, minimizing drastic adaptation that is sensitive to label shifts. Our selection scheme is based on two principles that we empirically motivate: (1) later layers of networks are more sensitive to label shift; (2) individual features can be sensitive to specific classes. We apply the proposed technique to three classification tasks, including CIFAR10-C, Imagenet-C, and diagnosis of fatty liver, where we explore both covariate and label distribution shifts. We find that our method allows to bring the benefits of TTA while significantly reducing the risk of failure common in imbalanced scenarios.
@@ 1 Introduction
Deep learning models have shown excellent performance on a wide variety of tasks in computer vision research, such as image classification (Krizhevsky et al., 2012; Yu et al., 2022), segmentation (Wasserthal et al., 2023; Ding et al., 2023), and object detection (Chen et al., 2020; Yin et al., 2022). However, a commonly cited limitation of these models is the inability to generalize across different domains (Hendrycks & Dietterich, 2019a). Generalization can be simply defined as the ability of a deep learning algorithm, trained on one or more source domains, to be applied to a different yet still related target domain. Typically, in real-world deployment scenarios models might encounter data with critical differences, hampering their performance. This decrease in performance has been observed in multiple areas, including life-threatening contexts, such as autonomous driving (Kim et al., 2022; VS et al., 2023) and medical diagnostics (Blaivas et al., 2022; Wang et al., 2021b).
@@
Domain adaptation techniques aim to enhance a model’s generalizability, anticipating the possible challenges in the event of distribution shifts. Traditionally, domain adaptation assumes that the model has access to unlabeled test data and source data during the training, which may not be practical in some cases (Piva et al., 2023; Liang et al., 2020). A recently emerging technique to deal with distribution shift is test-time adaptation (TTA) (Sun et al., 2020b; Liang et al., 2023), a type of unsupervised domain adaptation, where unlabeled test data is used to update the model parameters at test-time, before predictions. It is often assumed that data arrives in batches, and some studies have proposed a setting of test-time batch adaptation that take advantage of batch-level information to adapt to the distribution shift (Wang et al., 2021a; Nado et al., 2020; Chen et al., 2022).
@@
Test-time batch normalization (TTN) (Nado et al., 2020; Schneider et al., 2020) is an approach which replaces batch normalization statistics estimated as running averages on the training set with the statistics of the test data batch. Despite being a simple approach, it has been shown to improve robustness under covariate shift, handling particularly well various cases of image corruptions. Based on that, other TTA approaches apply TTN as a critical component in their foundation (Wang et al., 2021a; Lim et al., 2023; Zhao et al., 2023). However TTN was proposed and developed in the context of covariate shifts only, while in many realistic scenarios, such as medical imaging, the label distribution of data can shift from training to testing.
@@
In order to be practically useful, TTA methods relying on batch data need to be able to operate outside of this optimistic scenario of class-balanced training and testing settings, while simultaneously dealing with covariate shifts. As noted in recent work (Boudiaf et al., 2022) label distribution shifts can cause failures. We emphasize that these failures can be catastrophic, thus creating risk in the usage of TTA when the label distribution under deployment is unknown. We observe that certain channels and layers can be much more sensitive with respect to label distribution shifts under TTN. Motivated by this, we propose a method to mitigate the risk of label distribution shift catastrophically affecting model predictions. Our method, based on TTN, aims to adapt only those channels of the batch normalization layers whose activation statistics do not exhibit strong correlation to label distribution. The proposed method is applied to classification of two well-known benchmark natural image datasets — CIFAR-10 (Krizhevsky, 2012) and ImageNet-1K (Deng et al., 2009b) — and on adaptation from a private medical dataset (Vianna et al., 2023) to a publicly available one (Byra et al., 2018). When deployed in target data with different distribution, our proposed method is effective in obtaining the benefits of TTA while avoiding catastrophic failure cases, greatly improving the practical utility of TTN methods.
@@
    Experiments on three different datasets demonstrate that our method can be applied on a range of classification tasks, in particular for highly imbalanced target data. Notably, the adaptation across liver ultrasound datasets from different sites is shown to be effective. Our method is more robust than benchmark techniques across various degrees of label imbalance, showing similar performance in balanced cases and improvements in imbalanced cases.
@@ 2 Related Work
Batch normalization (Ioffe & Szegedy, 2015), henceforth BatchNorm, is a technique that aims to stabilize the distributions of layer inputs by introducing additional network layers that control the first two moments (mean and variance) of these distributions. This positively affects the optimization process of deep neural networks during training, reduces overfitting, and prevents the exploding or the vanishing gradient problems.
@@
Even though BatchNorm is effective, widely accepted, and commonly used in many deep neural networks, its mechanisms are not entirely understood. Methods extending this approach provide alternatives to BatchNorm, such as layer normalization (Ba et al., 2016), instance normalization (Ulyanov et al., 2016), and normalization based on groups of channels (Wu & He, 2018). Furthermore, weight normalization (Salimans & Kingma, 2016) extends this line of work, modifying the weights of the models instead of features.
@@
Various studies have been conducted to further understand how batch-normalization works. The original motivation for BatchNorm refers to the problem of internal covariate shift, however a later study (Santurkar et al., 2018) notes that it is not an appropriate explanation. Rather, the loss landscapes of models with BatchNorm appear to be smoother than those of equivalent models without it, allowing for use of larger range of learning rates and faster network convergence. Moreover, by investigating randomly initialized networks with all layers frozen and adapting only the BatchNorm layers, it has been shown that BatchNorm layers have substantial expressive power, achieving surprisingly good results on CIFAR-10 and ImageNet (Frankle et al., 2021).
@@
Test-time adaptation methods often involve updating the parameters associated with the BatchNorm layers to distributions with covariate shift (Schneider et al., 2020; Nado et al., 2020). This approach is closely related to the unsupervised domain adaptation method AdaBN (Li et al., 2016; Burns & Steinhardt, 2021), and has achieved promising results in mitigating image corruptions. Recent TTA methods have been published using entropy minimization on the BatchNorm affine parameters (Wang et al., 2021a), but they appear to be biased towards the dominant classes (Park et al., 2023). In contrast, a distinct class of methods focuses on TTA without relying on BatchNorm adjustments, proposing techniques that leverage ideas from self-supervised learning (Sun et al., 2020a; Chen et al., 2022), test-time augmentations (Hu et al., 2021; Khurana et al., 2021), or sample selection (Niu et al., 2022).
@@
A common limitation in TTA studies is regarding label distribution shifts, as most of the proposed methods fail to consider the varying class balance in the test sets, eventually falling into the “winner-takes-all” dilemma in which the majority classes dominate the predictions (You et al., 2021; Liang et al., 2023). To address the class-imbalanced domain adaptation challenge, some works focus on estimating the test set distribution (Sun et al., 2022; Lipton et al., 2018; Park et al., 2023).
@@
Recently, other works have also proposed methods for dealing with label distribution shift in TTA. LAME (Boudiaf et al., 2022) performs the online adaptation of the classification outputs instead of the model’s parameters, making it agnostic to training and testing conditions, but it does not perform well in balanced scenarios. DELTA (Zhao et al., 2023) uses test-time batch renormalization and dynamic online reweighting in an online manner to avoid mini-batch bias. A label shift adapter was designed (Park et al., 2023) to improve TTA performance in scenarios where both source and target are imbalanced, by estimating the distribution of the target set. However, it uses a large number of hyperparameters and is specialized to the case of imbalanced source data. Focusing on learned features, in few-shot learning it was observed that trained models presents channel bias, and a channel-wise transformation during test-time can improve predictions in out-of-distribution tasks (Luo et al., 2022).
@@ 3 Hybrid Test-Time Batch Normalization
In order to implement TTA in these settings, TTN views a neural network $n$ as split into blocks separated by BatchNorm layers:
$n=n_{0}\circ B_{0}^{s_{0}}\circ n_{1}\circ B_{1}^{s_{1}}\cdots n_{K-1}\circ B_{K-1}^{s_{K-1}}\circ n_{K},$ (1)
where $n_{0},\ldots,n_{K}$ are blocks (i.e., sub-networks) of hidden layers and $B_{0},\ldots,B_{K-1}$ are batch normalization operators. Each BatchNorm layer modifies each neural activation by
@@
Input: Trained network $n$ with $K$ layers; source model statistics $\{\bm{\mu}_{k}^{s},\bm{\sigma}_{k}^{s}\}_{k=1..K}$; source domain data $D$
for $c$ in $C$ classes do
       Sample Data from Class $c$
       $x_{0}\sim D_{c}$
       for $k$ in $\{1,...,K\}$ do
             $\bm{x}_{k}=n_{k-1}(\bm{x}_{k-1})$
             $\bm{\mu}_{k}^{t},\bm{\sigma}_{k}^{t}$leftarrow Compute BN stats from  $\bm{x}_{k}$
            for $f$ in $F$ channels do
                   Scores[c, k, f] = $W_{2}^{2}(\bm{\mu}_{f}^{s},\bm{\sigma}_{f}^{s},\bm{\mu}_{f}^{t},\bm{\sigma}_{f}^{t})$
Output:Scores
Algorithm 1 Hybrid-TTN Post-Training
for $k$ in $\{1,...,K\}$ do
       $\bm{x}_{k}=f_{k-1}(\bm{x}_{k-1})$
       $\bm{\mu}_{k}^{t},\bm{\sigma}_{k}^{t}$leftarrow Compute BN stats from  $\bm{x}_{k}$
       for $c$ in $C$ classes do
             for $f$ in $F$ channels do
                   Scoresk[f] += $p_{c}*$Scores[c, k, f]
      $R=\frac{k}{K}$
       topR-ind $\leftarrow$ Compute Top-R Index(Scoresk)
[…]
       $\bm{m}_{k}[\tilde{\textsc{topR-ind}}]=1$
       $\bm{\mu}_{k}^{hybrid}=\bm{m}_{k}\odot\bm{\mu}_{k}^{t}+(1-\bm{m}_{k})\odot\bm{\mu}_{k}^{s}$
       $\bm{\sigma}_{k}^{hybrid}=\bm{m}_{k}\odot\bm{\sigma}_{k}^{t}+(1-\bm{m}_{k})\odot\bm{\sigma}_{k}^{s}$
       $\bm{x}_{k}=B(\bm{x}_{k},\bm{\mu}_{k}^{hybrid},\bm{\sigma}_{k}^{hybrid})$
[…]
Algorithm 2 Hybrid-TTN Adaptation
@@
Storing Post-Training Statistics To mitigate the adverse effects by TTN, our first step focuses on selecting channels for the adaptation process. We aim to identify channels that are primarily sensitive to covariate shifts while excluding those highly sensitive to shifts in the label distribution. This channel selection procedure operates using an initial phase that uses the labelled training set before the model is deployed to compute a set of scores, completely independent of the target data . For each class, denoted as $c$ out of $C$ total classes, we input a subset of data from the class, $D_{c}$, into a trained neural network with $K$ layers. Each layer $k$ within this network possesses source statistics $\mu_{k}^{s}$ and $\sigma_{k}^{s}$, and the target statistics for the single class, $\mu_{k}^{t}$ and $\sigma_{k}^{t}$, are computed. The Wasserstein distance is then calculated for the source and target distributions with $W_{2}^{2}(\{\mu^{s},\sigma^{s}\},\{\mu^{t},\sigma^{t}\})=\|\mu^{s}-\mu^{t}\|^{2}+(\sigma^{s}-\sigma^{t})^{2}$. This step is described in Algorithm 1.
@@
At each layer we choose the top-$R$ channels for adaptation. Using the prior that later layers are more sensitive to label shifts, we decay $R$ with a simple linear function in depth. The algorithm is fully described in Algorithm 2.
In the method analyzed in our paper we focus on a simple method to obtain $p(c)$. Initially, Algorithm 2 is executed with a uniform prior to provide an initial corrected model. The network is prepared for adaptation using the uniform prior, the training data scores calculated in Algorithm 1, and the linear depth prior. By considering the uniform prior as weights to the individual class scores, selected channels in each layer are changed to use the target data batch statistics, with the remaining keeping the source model statistics. The model output is a first classification prediction, which can be further improved if we recalibrate the label distribution prior, using the first prediction.
Re-scoring with Pseudolabels Following the recomputation of the distribution prior using the first prediction, Algorithm 2 is rerun, incorporating the updated, non-uniform prior for further refinement, and giving the second round of predictions, which we consider our Hybrid-TTN output. This secondary application of Algorithm 2 can be seen as aiming to adjust the model prediction to exclude rare classes in the target distribution, unlike the first step which avoids updating channels sensitive to individual classes.